OpenStudy (romantic_ch3micals):

Anyone from Connexus and have answers for lesson 9 linear functions unit test algebra 1 9a?

OpenStudy (alexandervonhumboldt2):

post a question thats agains OS to ask for answers

OpenStudy (romantic_ch3micals):

...18. Write an equation for the line that is parallel to the given line and passes through the given point. y = 2x + 4; (3, 8)

OpenStudy (romantic_ch3micals):

.y = 2x + 2 y = 2x + 6 y = –2x + 6 y = negative 1/2x + 2

Directrix (directrix):

These two do: @MissTeacherLady16 @connectionsteacher

OpenStudy (alexandervonhumboldt2):

parsllel have the same slopes

OpenStudy (alexandervonhumboldt2):

1. y=2x+2

OpenStudy (alexandervonhumboldt2):

plz medal

OpenStudy (romantic_ch3micals):

-_- fine